OpenStudy (anonymous):

What are three different equivalent forms of x^2

OpenStudy (anonymous):

x^2 = x*x = (x^4)/(x^2)

OpenStudy (anonymous):

Also (x^-2)/(x^-4)

OpenStudy (anonymous):

It has to be a problem that can be solved with two or more steps. :(

OpenStudy (solomonzelman):

\[x^2=x \times x = x^{x} \div x^{x-2}=x^{1-x} \times x^{1 + x}\]

OpenStudy (solomonzelman):

and so forth

OpenStudy (anonymous):

(x+1)(x-1)+1

OpenStudy (anonymous):

Foil and you get x^2-1+1=x^2